OpenStudy (anonymous):

list the intercepts and test for symmetry y=x^3-512

OpenStudy (anonymous):

You could start by factoring. We have here the difference of two perfect cubes. The equation for this is: \[a^3 - b^3 = (a-b)(a^2 + ab +b^2)\] In this case, a=x and b=8. So: \[x^3 - 512 = (x-8)(x^2 +8a +64)\]

OpenStudy (anonymous):

so the intercepts are x=8 and y=-512

OpenStudy (anonymous):

Yep, you can graph it to check those or make sure there are no other roots.

OpenStudy (loser66):

@vinnv226 question: can I do as follow to find out the roots? x^3 -512 =0 --> x^3 = 512 x = cube root of 512 =8 by that way, I don't have any complex roots. Is it right?

OpenStudy (anonymous):

Thats definitely a valid way to find the roots. But there are two complex roots. You know this because it's a third degree, so it must have three roots, and there is only one real root.

OpenStudy (loser66):

Got you, thanks for explanation.

OpenStudy (loser66):

I still have another question. Would you mind explain me?

OpenStudy (anonymous):

If you're interested, the complex roots are 4(-1 - isqrt3) 4(-1 + isqrt3)

OpenStudy (loser66):

I 'll be back and make question later. I have a call. sorry

OpenStudy (anonymous):

thank you. @vinnv226